Solve. x2 – 75 = 0

Answer by Alan3354(69443)

Solve. x2 – 75 = 0
--------------
x^2 = 75
x = +sqrt(75) =~ 8.6603...
x = -sqrt(75) =~ -8.6603...
-------------
sqrt(75) can be written as 5sqrt(3)
Years ago, when people use log tables for square roots (among other things), it's was sometimes helpful to get the smallest number possible under the radical. That's not important with modern calculators, but some people, and some teachers, still fret about it.